Find an orthonormal basis of R^3, of which (0,3√13, 2√13) is one element

1
Expert's answer
2021-12-02T07:22:53-0500

The length of the vector "\vec v_1" is "1" as


"||\vec v_1 ||=\sqrt{(0)^2+(3/\sqrt{13})^2+(2/\sqrt{13})^2}=1"

We want to find two vectors "\vec v_2,\vec v_3" such that "\{\vec v_1,\vec v_2,\vec v_3\}"is an orthonormal basis for "\R^3."

Let "\vec u=\langle x, y, z \rangle" be a vector that is perpendicular to "\vec v_1."

Then we have "\vec u\cdot \vec v_1=0," and hence we have the relation


"0(x)+(3/\sqrt{13})(y)+(2/\sqrt{13})(z)=0"

"3y+2z=0"

For example, the vector "\vec u=\langle 1, 2, -3 \rangle" satisfies the relation, anf hence "\vec u\cdot \vec v_1=0."

Let us define the third vector "\vec w" to be the cross product of "\vec u" and "\vec v_1"


"\vec w=\vec u\times\vec v_1=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 1 & 2 & -3 \\
 0 & 3/\sqrt{13} & 2/\sqrt{13}
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 2 & -3 \\
 3/\sqrt{13} & 2/\sqrt{13}
\end{vmatrix}-\vec j\begin{vmatrix}
 1 & -3 \\
 0 & 2/\sqrt{13}
\end{vmatrix}"

"+\vec k\begin{vmatrix}
 1 & 2 \\
 0 & 3/\sqrt{13}
\end{vmatrix}"


"=(13/\sqrt{13})\vec i-(2/\sqrt{13})\vec j+(3/\sqrt{13})\vec k"

By the property of the cross product, the vector "\vec w" is perpendicular to both "\vec u, \vec v_1."

"||\vec u||=\sqrt{(1)^2+(2)^2+(-3)^2}=\sqrt{14}"

"||\vec w||=\sqrt{(13/\sqrt{13})^2+(-2/\sqrt{13})^2+(3/\sqrt{13})^2}"

"=\sqrt{14}"

Therefore, the set


"\{\vec v_1,\vec v_2,\vec v_3\}"


"=\bigg\{\begin{bmatrix}
 0 \\
 3/\sqrt{13}\\
2/\sqrt{13}
\end{bmatrix},\begin{bmatrix}
 1/\sqrt{14} \\
 2/\sqrt{14}\\
-3/\sqrt{14}
\end{bmatrix},\begin{bmatrix}
 13/\sqrt{182} \\
 3/\sqrt{182}\\
2/\sqrt{182}
\end{bmatrix}\bigg\}"

is an orthonormal basis for "\R^3."